When heated, lithium reacts with nitrogen to form lithium nitride:

               Li(s) + N2(g) ----------> Li3N(s)

What is the theoretical yield of Li3N in grams when 12.3 g of Li are heated with 33.6 g of N2?  

If the actual yield of Li3N is 5.89 g, what is the percent yield of the reaction?



1
Expert's answer
2021-01-27T04:36:50-0500

Balanced chemical reaction,

6Li(s) + N2(g) ---> 2Li3N

Moles of Li = mass/ molar mass

= 12.3/7.0 =1.76 mol

Moles of N2 = 33.6 /28 = 1.2 mol

Li is limiting reagent.

So, from balanced chemical reaction,

Molar mass of Li3N = 35 g

Mass of Li3N produced = 2×35×12.3/6×7

= 20.5 g

%yield = Actual yield×100/Theoretical yield

= 5.89×100/20.5

= 28.7 %